<e> Solve the first order linear inhomogeneous differential equation using the constant variation method

y,- (3y/x)=x



1
Expert's answer
2022-02-15T05:47:55-0500

Corresponding homogeneous differential equation


"y'-\dfrac{3y}{x}=0"

"\dfrac{dy}{y}=3\dfrac{dx}{x}"

Integrate


"\int \dfrac{dy}{y}=\int 3\dfrac{dx}{x}"

"\ln (|y|)=3\ln (|x|)+\ln C"

"y=Cx^3"

The general solution of the homogeneous differential equation is


"y_h=Cx^3"

Use the constant variation method


"y'=C'x^3+3x^2 C"

Substitute


"C'x^3+3x^2 C-\dfrac{3Cx^3}{x}=x"

"C'=x^{-2}"

Integrate


"C=\int x^{-2}dx"

"C=-\dfrac{1}{x}+C_1"

Finf the general solution of the non homogeneous differential equation


"y=(-\dfrac{1}{x}+C_1)x^3"

"y=-x^2+C_1x^3"